# Revealing Superconducting Chiral Edge Modes via Resistance Distributions

### v1

@@ Abstract
Edge transport with $N=1$ CBEM. Next, we consider the case where the middle SC region of the heterojunction has a BdG Chern number $N=1$, hosting a single self-conjugate CBEM, also known as chiral Majorana edge mode [6, 37]. As shown in Fig. 1(d), two incoherent incident charged modes, originating from different sources and labeled $a$ and $b$, decompose into four Majorana edge modes, labeled $\gamma_{1}$ through $\gamma_{4}$. Among them, $\gamma_{2}$ and $\gamma_{3}$ undergo braiding, after which all four modes recombine into two outgoing charged modes, labeled $c$ and $d$. This process, depicted in Fig. 1(e), is described by
@@
Here, $\Psi_{i}$ is a four-dimensional vector representing the initial state. The transformation sequence, proceeding from right to left, is as follows: $Q$ accounts for propagation near the corners upon entering the SC region, $M$ transforms the Nambu basis into the Majorana basis, $\Lambda$ describes propagation within the SC region in the Majorana basis, $S_{23}$ captures the braiding of $\gamma_{2}$ and $\gamma_{3}$, $M^{\prime}$ rotates the Majorana basis back to the Nambu basis, and $Q^{\prime}$ describes propagation near the corners before exiting the SC region. The matrix $\Lambda$ is diagonal, with random phase factors reflecting the well-separated and decoupled nature of the Majorana modes away from the SC region’s corners. However, near these corners, the Majorana modes overlap, leading to non-trivial coupling. This coupling takes the form $\eta\tau_{z}$ in the Nambu basis, where $\eta$ is a disorder-dependent coupling coefficient [40, 43]. This coupling term corresponds to the chemical potential term in $H_{\text{edge}}$. Consequently, the effect of random coupling is captured by $Q=\text{diag}(\mathbb{T}^{a},\mathbb{T}^{b})$ and $Q^{\prime}=\text{diag}(\mathbb{T}^{c},\mathbb{T}^{d})$, where $\mathbb{T}^{\zeta}$ represents the total transfer matrix at corner $\zeta=a,b,c,d$, structured similarly to the transfer matrix product in Eq. (2).
Case (iii): For the QAH-SC-QAH junction with $N=1$ CBEM shown in Fig. 1(d), $T$ consistently follows a normal distribution, regardless of the junction length.
The normal distribution of $T$ stems from two key factors. First, the incoherent nature of the incident electrons ensures that the phase factors in $\Lambda$ do not affect $T$, making it independent of the junction length. Second, coupling occurs only near the corners, leading to a small effective $L$ when computing $\mathbb{T}^{\zeta}$. As a result, each $\mathbb{T}^{\zeta}$ can be treated as a random rotation around a fixed axis, analogous to Case (i). A detailed derivation of this result is provided in [62].
@@
Then we examine an alternative junction configuration [24, 25, 27, 26, 28], depicted in Fig. 1(f). This setup, known as a QAH-SC junction, is realized by stacking a normal SC onto the edge region of a $C=1$ QAH insulator. When the SC region has a BdG Chern number $N=2$, the charge transmission fraction $T$ from lead 2 to 3 in QAH-SC junction matches that of the QAH-SC-QAH junction in Fig. 1(a). Therefore, we focus on the SC region has $N=1$. In this configuration, the incident charged mode $a$ splits into two Majorana modes, $\gamma_{1}$ and $\gamma_{2}$, within the SC region, which subsequently recombine and emerge as the outgoing charged mode $b$. A key distinction between this setup and the QAH-SC-QAH junction in Fig. 1(d) is that in the QAH-SC junction, the Majorana modes originate from the same source, preserving their coherence. This coherence leads to different transport behaviors of $T$, depending on the junction length. Specifically, the final state can be expressed as
@@
Here, fluctuations near the corners are negligible compared to those along the edges of the SC region. This allows us to approximate the total transfer matrix as $\overline{\mathbb{T}}^{-1}M^{\prime}\Lambda M\overline{\mathbb{T}}$, where $\overline{\mathbb{T}}$ represents the average transfer matrix at the corner. As a result, the entire process can be interpreted as a random rotation by an angle $\alpha$ around a fixed axis $\mathbf{n}$, with $\alpha$ and $\mathbf{n}$ determined by $\Lambda$ and $M\overline{\mathbb{T}}$, respectively. For small $L$, the fluctuation of $\alpha$ remains small, making the system behave similarly to Case (i). However, as $L$ increases, the system exhibits qualitatively different behavior.
Case (iv): For a long QAH-SC junction (large $L$ as $L>10^{4}$) with $N=1$ CBEM in Fig. 1(f), $T$ approximately follows a generalized arcsine distribution.
As discussed in Case (i), $\delta\alpha\propto\sqrt{L}$, which becomes significant for large $L$. Therefore, $\cos{\alpha}$ will converge to the standard arcsine distribution [67], then $T$ obeys a shifted and rescaled variant of the standard arcsine distribution, with the probability density function
@@
$f(T)=\begin{cases}\frac{1}{\pi\sqrt{(1-A)^{2}-(T-A)^{2}}}\ ,&(T\in[1-2A,1])\\ 0\ ,&(T\notin[1-2A,1])\end{cases}$ (5)
where $A=n_{z}^{2}$. This distribution is characterized by a U-shaped profile that peaks at the maximal value 1 and the minimal value ($1-2A$) of $T$.
Finally, when the SC region of the junction is topologically trivial ($N=0$), two CBEM emerge along the interface between the SC and QAH. Therefore, the edge mode transport characteristics closely resemble those of the $N=2$ case [62].
Numerical calculations. To validate the analytical results presented above, we numerically compute $T$ in each case. We implemented two-dimensional tight-binding Hamiltonians of both QAH and SC for the configurations in Fig. 1 (a,d,f) on a rectangular geometry. The disorder is introduced through a uniformly distributed onsite potential with $\xi_{d}=a\equiv 1$. The charge transmission $T$ is then calculated using the recursive Green’s function method [68, 69, 70, 71, 72, 73]. Additional numerical details are provided in [62].
@@
Figure 2: Mean and distribution of the charge transmission fraction $T$ for the junction in Fig. 1(a). (a,b) $T$ vs $L$, where solid lines and shaded region indicate the mean and standard deviation, respectively. (c-e) The distributions of $T$ for different $L$, with histograms representing numerical results and red lines showing fitted distributions. (c) The distribution is fitted with a normal distribution $f(T)=\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac{(T-E)^{2}}{2\sigma^{2}}}$. (d) The distribution is fitted with a chi-squared distribution $f(T)=\frac{1}{\sigma\sqrt{\pi(1-T)}}e^{-\frac{1-T}{\sigma^{2}}}$. (e) The mean and standard deviation of $T$ are $8.54\times 10^{-3}$ and 0.582, respectively, which match those of a uniform distribution $U[-1,1]$.
The results for the junction configuration of Fig. 1(a) are presented in Fig. 2. Here, $\overline{T}$ denotes the mean charge transmission fraction from lead 2 to 3. When disorder is introduced into the SC region, $\overline{T}$ exhibits oscillations with a decaying amplitude as the junction length $L$ increases, eventually converging to zero. Meanwhile, the fluctuations in $T$ increase, as shown in Fig. 2(a,b). For $L\gtrsim 2\ell_{d}$, $T$ follows a normal or chi-squared distribution, as illustrated in Fig. 2(c) and Fig. 2(d), respectively, in accordance with Case (i). For sufficiently large $L$, such that $L\gtrsim 10^{4}$, the distribution of $T$ approaches a continuous uniform distribution $U[-1,1]$, as shown in Fig. 2(e), corresponding to Case (ii).
@@
Figure 3: (a,b) Numerical results of $T$ for junction in Fig. 1(d). (a) $\overline{T}$ vs $L$, where $\varepsilon$ denotes the incident electron energy. The solid lines and shading indicate the mean and standard deviation, respectively. (b) Distribution of $T$ for $\varepsilon=0.28$, which follows a normal distribution depicted as red line. (c,d) Numerical results for junction of Fig. 1(f). (c) $\overline{T}$ vs $L$. (d) Distribution of $T$, with the red line showing the fitting distribution as a generalized arcsine distribution in Eq. (5).

### v2

@@ II.2 Edge transport with $N=1$ CBEM
Next, we consider the case where the middle SC region of the heterojunction has a BdG Chern number $N=1$, hosting a single self-conjugate CBEM, also known as chiral Majorana edge mode [6, 38]. As shown in Fig. 1(d), two incoherent incident charged modes, originating from different sources and labeled $a$ and $b$, decompose into four Majorana edge modes, labeled $\gamma_{1}$ through $\gamma_{4}$. Among them, $\gamma_{2}$ and $\gamma_{3}$ undergo braiding, after which all four modes recombine into two outgoing charged modes, labeled $c$ and $d$. This process, depicted in Fig. 1(e), is described by
@@
Here, $\Psi_{i}$ is a four-dimensional vector representing the initial state. The transformation sequence, proceeding from right to left, is as follows: $Q$ accounts for propagation near the corners upon entering the SC region. $M$ transforms the Nambu basis into the Majorana basis:
$M=\frac{1}{\sqrt{2}}\begin{pmatrix}1&1&0&0\\ -i&i&0&0\\ 0&0&1&1\\ 0&0&-i&i\end{pmatrix}$, (12)
under the basis $(\psi^{\dagger},\psi,\psi^{\prime\dagger},\psi^{\prime})^{\text{T}}$. $\Lambda$ describes propagation within the SC region in the Majorana basis. $S_{23}$ captures the braiding of $\gamma_{2}$ and $\gamma_{3}$:
$S_{23}=\begin{pmatrix}1&0&0&0\\ 0&0&1&0\\ 0&-1&0&0\\ 0&0&0&1\end{pmatrix}$. (13)
@@
$M^{\prime}=M^{-1}$ rotates the Majorana basis back to the Nambu basis. $Q^{\prime}$ describes propagation near the corners before exiting the SC region. The matrix $\Lambda$ is diagonal, whose diagonal elements are all phase factors, reflecting the well-separated and decoupled nature of the Majorana modes away from the SC region’s corners. However, near these corners, the Majorana modes overlap, leading to non-trivial coupling. This coupling takes the form $\eta\tau_{z}$ in the Nambu basis, where $\eta$ is a disorder-dependent coupling coefficient [41, 44]. This coupling term corresponds to the chemical potential term in $H_{\text{edge}}$. Consequently, the effect of random coupling is captured by $Q=\text{diag}(\mathbb{T}^{a},\mathbb{T}^{b})$ and $Q^{\prime}=\text{diag}(\mathbb{T}^{c},\mathbb{T}^{d})$, where $\mathbb{T}^{\zeta}=\begin{pmatrix}\zeta_{11}&\zeta_{12}\\ \zeta_{21}&\zeta_{22}\end{pmatrix}$ represents the total transfer matrix at corner $\zeta=a,b,c,d$, structured similarly to the transfer matrix product in Eq. (7).
As shown in Fig. 1(d), the two incoming electrons are phase-incoherent. The charge transmission from lead 2 to lead 3, denoted by $T$, is obtained by setting $\Psi_{i}=(1,0,0,0)^{\text{T}}$ and evaluating the first two components of $\Psi_{f}$. Similarly, the charge transmission from lead 5 to lead 3, denoted by $R$, is obtained by setting $\Psi_{i}=(0,0,1,0)^{\text{T}}$ and evaluating the first two components of $\Psi_{f}$. The final result is:
$T$ $=\frac{1}{4}|a_{11}+a_{21}|^{2}\left(|d_{11}+d_{12}|^{2}-|d_{21}+d_{22}|^{2}\right),$ (14)
$R$ $=\frac{1}{4}|b_{11}+b_{21}|^{2}\left(|d_{11}-d_{12}|^{2}-|d_{21}-d_{22}|^{2}\right).$
@@
Then we examine an alternative junction configuration [24, 25, 27, 26, 28], depicted in Fig. 1(f). This setup, known as a QAH-SC junction, is realized by stacking a normal SC onto the edge region of a $C=1$ QAH insulator. When the SC region has a BdG Chern number $N=2$, the charge transmission fraction $T$ from lead 2 to 3 in the QAH-SC junction matches that of the QAH-SC-QAH junction in Fig. 1(a). Therefore, we focus on the SC region that has $N=1$. In this configuration, the incident charged mode $a$ splits into two Majorana modes, $\gamma_{1}$ and $\gamma_{2}$, within the SC region, which subsequently recombine and emerge as the outgoing charged mode $b$. A key distinction between this setup and the QAH-SC-QAH junction in Fig. 1(d) is that in the QAH-SC junction, the Majorana modes originate from the same source, preserving their coherence. This coherence leads to different transport behaviors of $T$, depending on the junction length. Specifically, the final state can be expressed as
@@
Here, fluctuations near the corners are negligible compared to those along the edges of the SC region. This allows us to approximate the total transfer matrix as $\overline{\mathbb{T}}^{-1}M^{\prime}\Lambda M\overline{\mathbb{T}}$, where $\overline{\mathbb{T}}$ represents the average transfer matrix at the corner. As a result, the entire process can be interpreted as a random rotation by an angle $\alpha$ around a fixed axis $\mathbf{n}$, with $\alpha$ and $\mathbf{n}$ determined by $\Lambda$ and $M\overline{\mathbb{T}}$, respectively. The charge transmission fraction from lead 2 to 3 is the same as Eq. (10).
III Distribution of charge transmission fraction
The distribution of $T$ exhibits distinct yet general behaviors depending on the number of CBEM and the length of SC region, as analyzed in the following cases and summarized in Table 1.
III.1 Case (i)
For a relatively short QAH-SC-QAH junction ($L$ is small but $L>2\ell_{d}/a$, where $\ell_{d}$ is the decay length of the CBEM wavefuntion, and $a$ is the lattice constant) with $N=2$ CBEMs, as shown in Fig. 1(a), $T$ approximately follows either a normal distribution ($\mathcal{N}$), or a chi-squared distribution ($\chi^{2}$).
In this regime, fluctuations in the rotation axis are negligible, allowing us to approximate $\mathbf{n}\approx\overline{\mathbf{n}}_{\ell}$. In this case all $P_{\ell}$s become identical, denoted as $P_{1}$, so that $\Psi_{f}=P_{1}^{-1}\Lambda_{L}P_{1}\cdots P_{1}^{-1}\Lambda_{1}P_{1}\Psi_{i}=P_{1}^{-1}\Lambda_{L}\cdots\Lambda_{1}P_{1}\Psi_{i}$. As a result, the state undergoes sequential rotations around a nearly fixed axis, with each segment contributing a random rotation angle. The total rotation angle is $\alpha=\sum_{\ell=1}^{L}\alpha_{\ell}$. To exclude other effects such as crossed Andreev reflection and elastic cotunneling [67, 47, 68, 48], which would complicate the transport processes of interest, we require $L>2\ell_{d}/a$, which is typically at least of order 10. By the central limit theorem, after multiple segments, $\alpha$ follows a normal distribution with mean $\overline{\alpha}=L\overline{\alpha}_{\ell}$ and standard deviation $\delta\alpha=\sqrt{L}\delta\alpha_{\ell}$, irrespective of the specific distribution of the individual rotation angles $\alpha_{\ell}$. Because of the cosine term in Eq. (10), $T(\alpha)=A+(1-A)\cos{\alpha^{\prime}}$, where $A=n_{z}^{2},~\alpha^{\prime}=\alpha~\text{mod}~2\pi$. The normal distribution of $\alpha$ induces a wrapped normal distribution [69] of $\alpha^{\prime}$, with probability density function
$f_{\text{w}}(\alpha^{\prime})=\frac{1}{\sqrt{2\pi}\delta\alpha}\sum_{k=-\infty}^{+\infty}\exp\left[-\frac{(\alpha^{\prime}-\overline{\alpha^{\prime}}-2\pi k)^{2}}{2(\delta\alpha)^{2}}\right],$ (16)
where $\overline{\alpha^{\prime}}=\overline{\alpha}~\text{mod}~2\pi$. The probability density function of $T$ is then
$f(T)=\frac{f_{\text{w}}\left(\arccos{\left(\frac{T-A}{1-A}\right)}\right)+f_{\text{w}}\left(-\arccos{\left(\frac{T-A}{1-A}\right)}\right)}{(1-A)\sqrt{1-\left(\frac{T-A}{1-A}\right)^{2}}},$ (17)
for $T\in[2A-1,1]$ and zero otherwise. Since $\delta\alpha_{\ell}$ and $L$ are small, only one term in Eq. (16) contributes significantly, so $\alpha^{\prime}$ is approximately distributed as $N\left(\overline{\alpha^{\prime}},(\delta\alpha)^{2}\right)$. When $\overline{\alpha^{\prime}}$ is not close to 0 or $\pi$, expanding $\alpha^{\prime}$ around $\overline{\alpha^{\prime}}$ to first order, $\cos{\alpha^{\prime}}\approx\cos{\overline{\alpha^{\prime}}}-\sin{\overline{\alpha^{\prime}}}\Delta\alpha^{\prime}$, gives $T\approx A+(1-A)\cos{\overline{\alpha^{\prime}}}-(1-A)\sin{\overline{\alpha^{\prime}}}\Delta\alpha^{\prime}$, where $\Delta\alpha^{\prime}\sim N\left(0,(\delta\alpha)^{2}\right)$. Thus Eq. (17) reduces to
$f(T)\approx\frac{1}{\sqrt{2\pi}\sigma_{T}}e^{-\frac{(T-E)^{2}}{2\sigma_{T}^{2}}},$ (18)
where $\sigma_{T}=(1-A)\delta\alpha\sin\overline{\alpha^{\prime}}$ and $E=A+(1-A)\cos{\overline{\alpha^{\prime}}}$. Therefore, $T$ follows the normal distribution $N(E,\sigma_{T}^{2})$. When $\overline{\alpha^{\prime}}\approx 0$ or $\pi$, the first-order term vanishes and $\alpha^{\prime}$ must be expanded to the second order: $\cos{\alpha^{\prime}}\approx\pm 1\mp(\Delta\alpha^{\prime})^{2}/2$. Here $X=(\Delta\alpha^{\prime})^{2}$ follows a non-central chi-squared distribution with $k=1$, whose probability density function is: $f(X)=e^{-\frac{X}{2(\delta\alpha)^{2}}}/(\delta\alpha\sqrt{2\pi X})$. Thus the probability density function of $T\approx A+(1-A)(\pm 1\mp X/2)$ is:
@@ III.1 Case (i)
$f(T)=\begin{cases}\frac{1}{\sqrt{\pi(1-T)}\sigma_{T}}e^{-\frac{1-T}{\sigma_{T}^{2}}},&\overline{\alpha^{\prime}}=0\\ \frac{1}{\sqrt{\pi(T-(2A-1))}\sigma_{T}}e^{-\frac{T-(2A-1)}{\sigma_{T}^{2}}},&\overline{\alpha^{\prime}}=\pi\\ \end{cases},$ (19)
where $\sigma_{T}=\sqrt{1-A}\delta\alpha$. This implies that $T$ follows a chi-squared distribution. As the junction length varies, the distribution of $T$ oscillates between normal and chi-squared distributions. This behavior extends beyond the results in Ref. [49], where only a normal distribution was proposed.
III.2 Case (ii)
For a long QAH-SC-QAH junction with (sufficiently large $L$ and) $N=2$ CBEM, as shown in Fig. 1(a), $T$ approximates a continous uniform distribution over [1,1][-1,1], denoted as U[1,1]U[-1,1].
As $L$ increases, the fluctuations of the rotation axis become significant, causing the state to spread across the entire Bloch sphere, as illustrated in Fig. 1(c). The uniform distribution of transmission can be explained using the theorem of Itô and Kawada [70]: Consider a stochastic process on a group $G$: $S_{L}=R_{1}R_{2}\cdots R_{L}$, where $R_{i}$ are independent random variables distributed according to a single probability measure $\mu$ on $G$. The probability measure of $S_{L}$, denoted $\mu^{L}$, is the $L$-fold convolution of $\mu$. If $\mu$ is aperiodic, then the sequence $\mu^{L}$ converges in distribution to the normalized Haar measure on $G$.
Here we focus on the group $G=SU(2)$. The fluctuation region of $\mathbf{n}_{\ell}$ defines the probability measure $\mu$, whose support is not contained in a coset of any proper closed subgroup of $G$, ensuring aperiodicity. A sequence of such random rotations thus forms a stochastic process on $G$. According to the theorem, after sufficiently many rotations, the probability measure converges to the normalized Haar measure of $SU(2)$, implying that $\Psi_{f}$ becomes equidistributed on the Bloch sphere. Consequently, the $z$ component of $\Psi_{f}$ follows U[1,1]U[-1,1]. Parametrizing the wavefunction as $\Psi_{f}=(\cos(\theta/2),e^{i\phi}\sin(\theta/2))^{\text{T}}$, we obtain $T=\cos{\theta}=z\sim U[-1,1]$, i.e., a continuous uniform distribution.
III.3 Case (iii)
For the QAH-SC-QAH junction with $N=1$ CBEM shown in Fig. 1(d), $T$ consistently follows a normal distribution, regardless of the junction length.
@@ III.3 Case (iii)
Figure 2: Mean and distribution of the charge transmission fraction $T$ for the junction in Fig. 1(a). (a), (b) $\overline{T}$ vs $L$, where solid lines and shaded region indicate the mean and standard deviation, respectively. (c-e) The distributions of $T$ for different $L$, with histograms representing numerical results and red lines showing fitted distributions. (c) The distribution is fitted with a normal distribution $f(T)=\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac{(T-E)^{2}}{2\sigma^{2}}}$. (d) The distribution is fitted with a chi-squared distribution $f(T)=\frac{1}{\sigma\sqrt{\pi(1-T)}}e^{-\frac{1-T}{\sigma^{2}}}$. (e) The mean and standard deviation of $T$ are $8.54\times 10^{-3}$ and 0.582, respectively, which match those of a uniform distribution U[1,1]U[-1,1].
In this case, the total transfer matrix at corner $\zeta^{a}$ can also be parametrized as $\mathbb{T}^{a}=e^{i\alpha(\mathbf{n}\cdot\bm{\sigma})/2}$, thus:
$|a_{11}\pm a_{21}|^{2}=1\pm n_{x}n_{y}-\mp n_{x}n_{z}\cos{\alpha}\mp n_{y}\sin{\alpha}.$ (20)
Because the edge-mode wavefunction has a finite decay length, $\eta\neq 0$ only near the corner. As a result, the effective $L$ is small, resembling case (i). Expanding around $\overline{a}$ to first order, $|a_{11}+a_{21}|^{2}\sim\overline{a}+\Delta a$ is normally distributed with small variance. Similar results hold for $|a_{11}-a_{21}|^{2}$, $|b_{11}\pm b_{21}|^{2}$, and related terms. Consequently, $|a_{11}+a_{21}|^{2}|d_{11}+d_{12}|^{2}\sim\overline{a}\overline{d}+\overline{a}\Delta d+\overline{d}\Delta a$ is also normally distributed to first order. Therefore, both $T$ and $R$ follow normal distributions with small variance. The above analysis reveals that the transmission probability is independent of $L$, which results from the incoherence of incident electrons, making the phase accumulation in $\Lambda$ in Eq. (11) have no effect on the final result.
III.4 Case (iv)
For a long QAH-SC junction (sufficiently large $L$) with $N=1$ CBEM, as shown in Fig. 1(f), $T$ approximately follows a generalized arcsine distribution.
As discussed in case (i), $\delta\alpha\propto\sqrt{L}$, which becomes significant for large $L$. Therefore, $\cos{\alpha}$ will converge to the standard arcsine distribution [71]. This can be seen by expanding Eq. (16) in a Fourier series:
$f_{\text{w}}(\alpha^{\prime})=\frac{1}{2\pi}\sum_{n=-\infty}^{\infty}e^{in(\alpha^{\prime}-\overline{\alpha^{\prime}})-\frac{(\delta\alpha)^{2}n^{2}}{2}}.$ (21)
When $\delta\alpha\rightarrow\infty$, only the $n=0$ term contributes, and $f_{\text{w}}(\alpha^{\prime})$ reduces to a constant $1/(2\pi)$, which implies that $\alpha^{\prime}$ is uniformly distributed as $U[0,2\pi]$. Then Eq. (17) approaches
$f(T)=\begin{cases}\frac{1}{\pi\sqrt{(1-A)^{2}-(T-A)^{2}}}\ ,&(T\in[2A-1,1])\\ 0\ ,&(T\notin[2A-1,1])\end{cases}$ (22)
Thus, $T$ follows a shifted and rescaled variant of the standard arcsine distribution, characterized by a U-shaped profile that diverges near the endpoints at $T=1$ and T=2A1T=2A-1.
@@ III.4 Case (iv)
Figure 3: (a), (b) Numerical results of $T$ for junction in Fig. 1(d). (a) $\overline{T}$ vs $L$, where $\varepsilon$ denotes the incident electron energy. The solid lines and shading indicate the mean and standard deviation, respectively. (b) Distribution of $T$ for $\varepsilon=0.28$, which follows a normal distribution depicted as a red line. (c), (d) Numerical results for junction of Fig. 1(f). (c) $\overline{T}$ vs $L$. (d) Distribution of $T$, with the red line showing the fitting distribution as a generalized arcsine distribution in Eq. (22).
Finally, we make the following remarks:
(a) When the SC region of the junction is topologically trivial ($N=0$), two CBEMs appear along the SC-QAH interface. The transport properties in this case closely resemble those of the $N=2$ junction [64].
(b) There is no universal threshold for a “sufficiently large” $L$, as it depends on disorder, incident energy, and other parameters of the system. For smaller incident energies, a longer junction is required to reach convergence in the distribution of $T$. In contrast, a stronger disorder leads to convergence at shorter lengths.
(c) The explanation of case (iii) is based on the central limit theorem, which assumes a sufficiently large number of random rotations. Deviations may occur when the edge-mode decay length is short or the disorder correlation length is long, both of which reduce the effective number of independent rotations, potentially altering the distribution of $T$.
IV Numerical simulations
To validate the analytical results presented above, we numerically compute $T$ in each case. We implemented two-dimensional tight-binding Hamiltonians of both QAH and SC for the configurations in Figs. 1 (a), 1(d), and 1(f) on a rectangular geometry. The disorder is introduced through a uniformly distributed onsite potential and spatially uncorrelated. We set $a\equiv 1$. The charge transmission $T$ is then calculated using the recursive Green’s function method [72, 73, 74, 75, 76, 77]. Additional numerical details are provided in Ref. [64].
The results for the junction configuration of Fig. 1(a) are presented in Fig. 2. Here, $\overline{T}$ denotes the mean charge transmission fraction from lead 2 to 3. When disorder is introduced into the SC region, $\overline{T}$ exhibits oscillations with a decaying amplitude as the junction length $L$ increases, eventually converging to zero. Meanwhile, the fluctuations in $T$ increase, as shown in Figs. 2(a) and 2(b). For $L\gtrsim 2\ell_{d}$, $T$ follows a normal or chi-squared distribution, as illustrated in Fig. 2(c) and Fig. 2(d), respectively, in accordance with case (i). For sufficiently large $L$, in this parameter setup $L\gtrsim 10^{4}$, the distribution of $T$ approaches a continuous uniform distribution U[1,1]U[-1,1], as shown in Fig. 2(e), corresponding to case (ii).